OpenStudy (anonymous):

A cube has a volume of 5 cm3. Each edge is tripled in length. What is the volume of the new cube?

OpenStudy (ja1):

15, also love the name XD

OpenStudy (ja1):

@JessicaDoesntLikeMath

OpenStudy (anonymous):

\[x^3=5 \]\[x=\sqrt[3]{5}\]

OpenStudy (anonymous):

\[\left(3\sqrt[3]{5}\right)^3=135 \]